OpenStudy (anonymous):

Find the area of the shaded region. a. 72 square units c. 252 square units b. 45 square units d. 288 square units

Mehek (mehek14):

attachment?

OpenStudy (anonymous):



OpenStudy (anonymous):

the area= x^2

OpenStudy (anonymous):

idk what to do

OpenStudy (anonymous):

x^2=3^2+6^2=9+36=45

OpenStudy (luigi0210):

You sure about that @AJ01 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the shaded region is square the square area= L^2 so you need the length of L from the plot L is hyp of right triangle

OpenStudy (luigi0210):

I think it would go something like this: Since all sides are the same, it is a square: Adding the sides, 9. So the area of the square is 9^2=81 To find the shaded, you need to subtract the unshaded, which are triangles with a height of 6 and base of 3, and there are 4: \(\Large 9^2-4(\frac{1}{2}*6*3) \) \(\Large 81-4(9) \) \(\Large 81-36 \) And yesss